Three identical dice are rolled. The  probability that distinct number appears on them is
 

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Kshitiz
 
Distinct number should be on each one of them,
Then total number of ways = 6*5*4 = 120
Total possible outcomes = 6*6*6 = 216
 
Hence probability = 120/216 = 20/36 = 5/9
 
Regards
Arun (askIITians forum expert)